# Finite-Sample Analysis of Policy Evaluation for Robust Average Reward Reinforcement Learning

### v1

@@ Appendix B Biased Semi-Norm Stochastic Approximation
Note that since $\|\cdot\|_{s,\overline{E}}$ and $\|\cdot\|_{\infty,\overline{E}}$ are semi-norms on a finite‐dimensional space with the same kernel, there exist $\rho_{1},\rho_{2}>0$ such that
$\rho_{1}\,\|z\|_{\infty,\overline{E}}\leq\|z\|_{s,\overline{E}}\leq\rho_{2}\,\|z\|_{\infty,\overline{E}},\;\forall\,z.$ (55)
Likewise, their dual norms (denoted $\|\cdot\|_{*,s,\overline{E}}$ and $\|\cdot\|_{*,\infty,\overline{E}}$) satisfy the following:
$\frac{1}{\rho_{2}}\|z\|_{*,s,\overline{E}}\leq\|z\|_{*,\infty,\overline{E}}\leq\frac{1}{\rho_{1}}\|z\|_{*,s,\overline{E}},\;\forall\,z.$ (56)
B.3 Formal Statement of Theorem 3
@@
In the biased noise setting, $u=\nabla M_{\overline{E}}(x^{t}-x^{*})$ and $v=\mathbb{E}[w^{t}|\mathcal{F}^{t}]$, with $\|\cdot\|=\|\cdot\|_{\infty,\overline{E}}$. So
$\bigl{\langle}\nabla M_{\overline{E}}(x^{t}-x^{*}),\mathbb{E}[w^{t}|\mathcal{F}^{t}]\bigr{\rangle}\leq\bigl{\|}\nabla M_{\overline{E}}(x^{t}-x^{*})\bigr{\|}_{*,\,\infty,\overline{E}}\cdot\bigl{\|}\mathbb{E}[w^{t}|\mathcal{F}^{t}]\bigr{\|}_{\infty,\overline{E}}.$ (60)
Since $\|\mathbb{E}[w^{t}|\mathcal{F}^{t}]\|_{\infty,\overline{E}}\leq\varepsilon_{\text{bias}}$, it remains to bound $\|\nabla M_{\overline{E}}(x^{t}-x^{*})\|_{*,\infty,\overline{E}}$. By setting $y=0$ in (54), we get
@@
$\|\nabla M_{\overline{E}}(x)\|_{*,\infty,\overline{E}}\leq\alpha\|\nabla M_{\overline{E}}(x)\|_{*,s,\overline{E}}$ (63)
@@
$\|\nabla M_{\overline{E}}(x)\|_{*,\infty,\overline{E}}\leq\alpha\big{(}\|\nabla M_{\overline{E}}(0)\|_{*,s,\overline{E}}+L\|x\|_{s,\overline{E}}\big{)},\quad\forall\,x.$ (64)
By (55), we know that $\|x\|_{s,\overline{E}}\leq\|x\|_{\infty,\overline{E}}$, thus we have:
$\|\nabla M_{\overline{E}}(x)\|_{*,\infty,\overline{E}}\leq\alpha\big{(}\|\nabla M_{\overline{E}}(0)\|_{*,s,\overline{E}}+L\rho_{2}\|x\|_{\infty,\overline{E}}\big{)},\quad\forall\,x.$ (65)
@@
$\mathbb{E}\Bigl{[}\langle\nabla M_{\overline{E}}(x^{t}-x^{*}),\,w^{t}\rangle|\mathcal{F}^{t}\Bigr{]}=\bigl{\langle}\nabla M_{\overline{E}}(x^{t}-x^{*}),\mathbb{E}[w^{t}|\mathcal{F}^{t}]\bigr{\rangle}\leq G\bigl{(}1+\|x^{t}-x^{*}\|_{\infty,\overline{E}}\bigr{)}\varepsilon_{\text{bias}}.$ (67)
Combining (67) with (B.3) we obtain
$\mathbb{E}[\langle\nabla M_{\overline{E}}(x^{t}-x^{*}),x^{t+1}-x^{t}\rangle|\mathcal{F}^{t}]\leq\eta_{t}\langle\nabla M_{\overline{E}}(x^{t}-x^{*}),H(x^{t})-x^{t}\rangle+\eta_{t}G\varepsilon_{\text{bias}}\Bigl{(}1+\|x^{t}-x^{*}\|_{\infty,\overline{E}}\Bigr{)}$ (68)
To bound the first term in the RHS of (68), note that
@@
where $(a)$ follows from the convexity of $M_{\bar{E}}$, $(b)$ follows from $x^{*}$ belonging to a fixed equivalent class with respect to $H$ and $(c)$ follows from the contraction property of $H$. Combining (B.3). (68) and Lemma 4 with (57), we arrive as follows:
@@
$+\eta_{t}G\varepsilon_{\text{bias}}\Bigl{(}1+\|x^{t}-x^{*}\|_{\infty,\overline{E}}\Bigr{)}$ (70)
@@
let $\alpha_{2}$, $\alpha_{3}$ and $\alpha_{4}$ be defined in (B.3), if $x^{t}$ is generated by (51) with all assumptions in B.1 satisfied, then if the stepsize $\eta_{t}\coloneqq\frac{1}{\alpha_{2}(t+K)}$ while $K\coloneqq\max\{\alpha_{3}/\alpha_{2},3\}$,
$\mathbb{E}\Bigl{[}\|x^{T}-x^{*}\|^{2}_{\infty,\overline{E}}\Bigr{]}\leq\frac{K^{2}c_{u}}{(T+K)^{2}c_{l}}\|x^{0}-x^{*}\|^{2}_{\infty,\overline{E}}+\frac{8\alpha_{4}c_{u}}{(T+K)\alpha_{2}^{2}}+\frac{2c_{u}C_{1}C_{2}\varepsilon_{\text{bias}}}{\alpha_{2}}$ (71)
where $C_{1}=G(1+2x_{sp})$, $C_{2}=\frac{1}{K}+\log\big{(}\frac{T-1+K}{K}\big{)}$, $G$ is defined in (66) and $x_{sp}\coloneqq\sup\|x\|_{\infty,\overline{E}}$ is the upper bound of the span for all $x^{t}$.
@@
This choice $\eta_{t}$ satisfies $\alpha_{3}\eta_{t}^{2}\leq\alpha_{2}\eta_{t}$. Thus, by (B.3) we have
@@
where the term $R_{1}$ is identical to the unbiased case in Theorem 3 of (Zhang et al., 2021) which leads to
@@
also, $R_{2}$ can be bounded by a logrithmic dependence of $T$
@@
Appendix C Uncertainty Set Support Function Estimators
C.1 Proof of Theorem 4
@@
C.2 Proof of Theorem 5
denote $\hat{\sigma}^{*}_{\mathcal{P}^{a}_{s}}(V)$ as the untuncated MLMC estimator obtained by running Algorithm 1 when setting $N_{\max}$ to infinity. From (Wang et al., 2023b), under both TV uncertainty sets and Wasserstein uncertainty sets, we have $\hat{\sigma}^{*}_{\mathcal{P}^{a}_{s}}(V)$ as an unbiased estimator of ${\sigma}_{\mathcal{P}^{a}_{s}}(V)$. Thus,
@@ C.3 Proof of Lemma 2
Combining (92) with (C.3) we obtain (28).
@@
using (96) and the fact that $|f(\lambda)-g(\lambda)|\leq\epsilon\Rightarrow|\sup_{\lambda}f(\lambda)-\sup_{\lambda}g(\lambda)|\leq\epsilon$, we obtain (29).
C.4 Proof of Theorem 6
@@
Appendix D Convergence for Robust TD
D.1 Formal Statement of Theorem 7
The first half of Algorithm 2 (line 1 - line 7) can be treated as a special instance of the SA updates in (51) with the bias and variance of the i.i.d. noise term specified in Section 6. We start with analyzing the bias and variance of $\hat{\mathbf{T}}_{g_{0}}(V_{t})$ for each $t$. Recall the definition of $\hat{\mathbf{T}}_{g_{0}}(V_{t})$ is as follows:
@@ D.1 Formal Statement of Theorem 7
$\mathbb{E}\left[(\hat{\mathbf{T}}_{g_{0}}(V_{t})(s)-{\mathbf{T}}_{g_{0}}(V_{t})(s))^{2}\right]$ $=\left(\mathbb{E}\left[\hat{\mathbf{T}}_{g_{0}}(V_{t})(s)\right]-{\mathbf{T}}_{g_{0}}(V_{t})(s)\right)^{2}+\mathrm{Var}\left(\hat{\mathbf{T}}_{g_{0}}(V_{t})(s)\right)$
@@
$t^{p}_{\mathrm{mix}}<+\infty\quad\text{and}\quad\|V\|_{\mathrm{sp}}\leq 4t^{p}_{\mathrm{mix}}$ (105)
Thus, define $t_{\mathrm{mix}}\coloneqq\sup_{p\in\mathcal{P}}t^{p}_{\mathrm{mix}}$, then $t_{\mathrm{mix}}$ is also finite due to the compactness of $\mathcal{P}$. We now derive the bounds of biases and variances for the three types of uncertainty sets. Regarding contamination uncertainty sets, according to Lemma 5, $\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)$ is unbiased and has variance bounded by $\|V\|^{2}$. Thu, define $t_{\mathrm{mix}}$ according to (142) and combining the above result with Lemma 6, we obtian that $\hat{\mathbf{T}}_{g_{0}}(V_{t})$ is also unbiased and the variance satisfies
@@
$\mathbb{E}\left[\left\lVert\hat{\mathbf{T}}_{g_{0}}(V_{t})-{\mathbf{T}}_{g_{0}}(V_{t})\right\rVert^{2}_{\infty}\right]$ $\leq\left(24(1+\frac{1}{\delta})2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}\right)^{2}+3\|V\|_{\mathrm{sp}}^{2}+144(1+\frac{1}{\delta})^{2}\|V\|_{\mathrm{sp}}^{2}N_{\max}$
@@
Thus, using the fact that the span is less or equal to two times the $l_{\infty}$ norm, line 1 - line 7 of Algorithm 2 can be treated as a special instance of the SA updates in (51) with $B=0$. Furthermore, for contamination uncertainty sets, we have
$\varepsilon^{\text{Cont}}_{\text{bias}}=0\quad\text{and}\quad A^{\text{Cont}}=16t^{2}_{\mathrm{mix}}$ (111)
@@
$\varepsilon^{\text{TV}}_{\text{bias}}=48(1+\frac{1}{\delta})2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}=\operatorname*{\mathcal{O}}\left(2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}\right)$ (112)
$A^{\text{TV}}=2\left(24(1+\frac{1}{\delta})2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}\right)^{2}+96t_{\mathrm{mix}}^{2}+4608(1+\frac{1}{\delta})^{2}t_{\mathrm{mix}}^{2}N_{\max}=\operatorname*{\mathcal{O}}\left(t_{\mathrm{mix}}^{2}N_{\max}\right)$ (113)
@@
$\varepsilon^{\text{Wass}}_{\text{bias}}=48\cdot 2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}=\operatorname*{\mathcal{O}}\left(2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}\right)$ (114)
$A^{\text{Wass}}=2\left(24\cdot 2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}\right)^{2}+96t_{\mathrm{mix}}^{2}+4608t_{\mathrm{mix}}^{2}N_{\max}=\operatorname*{\mathcal{O}}\left(t_{\mathrm{mix}}^{2}N_{\max}\right)$ (115)
@@
$\mathbb{E}\Bigl{[}\|V_{T}-V^{*}\|^{2}_{\mathrm{sp}}\Bigr{]}\leq\frac{K^{2}c_{u}}{(T+K)^{2}c_{l}}\|V_{0}-V^{*}\|^{2}_{\mathrm{sp}}+\frac{4A^{\mathrm{Cont}}\alpha_{4}c_{u}}{(T+K)\alpha_{2}^{2}}=\operatorname*{\mathcal{O}}\left(\frac{1}{T^{2}}+\frac{t^{2}_{\mathrm{mix}}}{T(1-\gamma)^{2}}\right)$ (116)
@@
$\mathbb{E}\Bigl{[}\|V_{T}-V^{*}\|^{2}_{\mathrm{sp}}\Bigr{]}$ $\leq\frac{K^{2}c_{u}}{(T+K)^{2}c_{l}}\|V_{0}-V^{*}\|^{2}_{\mathrm{sp}}+\frac{4A^{\mathrm{TV}}\alpha_{4}c_{u}}{(T+K)\alpha_{2}^{2}}+\frac{c_{u}C_{3}C_{2}\varepsilon^{\mathrm{TV}}_{\mathrm{bias}}}{\alpha_{2}}$ (117)
@@
$\mathbb{E}\Bigl{[}\|V_{T}-V^{*}\|^{2}_{\mathrm{sp}}\Bigr{]}$ $\leq\frac{K^{2}c_{u}}{(T+K)^{2}c_{l}}\|V_{0}-V^{*}\|^{2}_{\mathrm{sp}}+\frac{4A^{\mathrm{Wass}}\alpha_{4}c_{u}}{(T+K)\alpha_{2}^{2}}+\frac{c_{u}C_{3}C_{2}\varepsilon^{\mathrm{Wass}}_{\mathrm{bias}}}{\alpha_{2}}$ (119)
@@
where $C_{2}=\frac{1}{K}+\log\big{(}\frac{T-1+K}{K}\big{)}$, $C_{3}=G(1+8t_{\mathrm{mix}})$, $\gamma$ is defined in (17), $c_{u},c_{l}$ are defined in (52), $\rho_{2}$ is defined in (55) and $G$ is defined in (66).

### v2

@@ D.1 Proof of Theorem 8
Note that since $\|\cdot\|_{s,\overline{E}}$ and $\|\cdot\|_{\mathcal{N},\overline{E}}$ are semi-norms on a finite‐dimensional space with the same kernel, there exist $\rho_{1},\rho_{2}>0$ such that
$\rho_{1}\,\|z\|_{\mathcal{N},\overline{E}}\leq\|z\|_{s,\overline{E}}\leq\rho_{2}\,\|z\|_{\mathcal{N},\overline{E}},\;\forall\,z.$ (70)
Likewise, their dual norms (denoted $\|\cdot\|_{*,s,\overline{E}}$ and $\|\cdot\|_{*,\mathcal{N},\overline{E}}$) satisfy the following:
$\frac{1}{\rho_{2}}\|z\|_{*,s,\overline{E}}\leq\|z\|_{*,\mathcal{N},\overline{E}}\leq\frac{1}{\rho_{1}}\|z\|_{*,s,\overline{E}},\;\forall\,z.$ (71)
D.1.3 Formal Statement of Theorem 8
@@
In the biased noise setting, $u=\nabla M_{\overline{E}}(x^{t}-x^{*})$ and $v=\mathbb{E}[w^{t}|\mathcal{F}^{t}]$, with $\|\cdot\|=\|\cdot\|_{\mathcal{N},\overline{E}}$. So
$\bigl{\langle}\nabla M_{\overline{E}}(x^{t}-x^{*}),\mathbb{E}[w^{t}|\mathcal{F}^{t}]\bigr{\rangle}\leq\bigl{\|}\nabla M_{\overline{E}}(x^{t}-x^{*})\bigr{\|}_{*,\,\mathcal{N},\overline{E}}\cdot\bigl{\|}\mathbb{E}[w^{t}|\mathcal{F}^{t}]\bigr{\|}_{\mathcal{N},\overline{E}}.$ (75)
Since $\|\mathbb{E}[w^{t}|\mathcal{F}^{t}]\|_{\mathcal{N},\overline{E}}\leq\varepsilon_{\text{bias}}$, it remains to bound $\|\nabla M_{\overline{E}}(x^{t}-x^{*})\|_{*,\mathcal{N},\overline{E}}$. By setting $y=0$ in (69), we get
@@
$\|\nabla M_{\overline{E}}(x)\|_{*,\mathcal{N},\overline{E}}\leq\alpha\|\nabla M_{\overline{E}}(x)\|_{*,s,\overline{E}}$ (78)
@@
$\|\nabla M_{\overline{E}}(x)\|_{*,\mathcal{N},\overline{E}}\leq\alpha\big{(}\|\nabla M_{\overline{E}}(0)\|_{*,s,\overline{E}}+L\|x\|_{s,\overline{E}}\big{)},\quad\forall\,x.$ (79)
By (70), we know that $\|x\|_{s,\overline{E}}\leq\|x\|_{\mathcal{N},\overline{E}}$, thus we have:
$\|\nabla M_{\overline{E}}(x)\|_{*,\mathcal{N},\overline{E}}\leq\alpha\big{(}\|\nabla M_{\overline{E}}(0)\|_{*,s,\overline{E}}+L\rho_{2}\|x\|_{\mathcal{N},\overline{E}}\big{)},\quad\forall\,x.$ (80)
@@
$\mathbb{E}\Bigl{[}\langle\nabla M_{\overline{E}}(x^{t}-x^{*}),\,w^{t}\rangle|\mathcal{F}^{t}\Bigr{]}=\bigl{\langle}\nabla M_{\overline{E}}(x^{t}-x^{*}),\mathbb{E}[w^{t}|\mathcal{F}^{t}]\bigr{\rangle}\leq G\bigl{(}1+\|x^{t}-x^{*}\|_{\mathcal{N},\overline{E}}\bigr{)}\varepsilon_{\text{bias}}.$ (82)
Combining (82) with (D.1.3) we obtain
$\mathbb{E}[\langle\nabla M_{\overline{E}}(x^{t}-x^{*}),x^{t+1}-x^{t}\rangle|\mathcal{F}^{t}]\leq\eta_{t}$ $\langle\nabla M_{\overline{E}}(x^{t}-x^{*}),H(x^{t})-x^{t}\rangle$
$+\eta_{t}G\varepsilon_{\text{bias}}\Bigl{(}1+\|x^{t}-x^{*}\|_{\mathcal{N},\overline{E}}\Bigr{)}$ (83)
To bound the first term in the RHS of (D.1.3), note that
@@
where $(a)$ follows from the convexity of $M_{\bar{E}}$, $(b)$ follows from $x^{*}$ belonging to a fixed equivalent class with respect to $H$ and $(c)$ follows from the contraction property of $H$. Combining (D.1.3). (D.1.3) and Lemma 11 with (72), we arrive as follows:
@@
$+\eta_{t}G\varepsilon_{\text{bias}}\Bigl{(}1+\|x^{t}-x^{*}\|_{\mathcal{N},\overline{E}}\Bigr{)}$ (85)
@@
let $\alpha_{2}$, $\alpha_{3}$ and $\alpha_{4}$ be defined in (D.1.3), if $x^{t}$ is generated by (66) with all assumptions in D.1.1 satisfied, then if the stepsize $\eta_{t}\coloneqq\frac{1}{\alpha_{2}(t+K)}$ while $K\coloneqq\max\{\alpha_{3}/\alpha_{2},3\}$,
$\mathbb{E}\Bigl{[}\|x^{T}-x^{*}\|^{2}_{\mathcal{N},\overline{E}}\Bigr{]}\leq\frac{K^{2}c_{u}}{(T+K)^{2}c_{l}}\|x^{0}-x^{*}\|^{2}_{\mathcal{N},\overline{E}}+\frac{8\alpha_{4}c_{u}}{(T+K)\alpha_{2}^{2}}+\frac{2c_{u}C_{1}C_{2}\varepsilon_{\text{bias}}}{\alpha_{2}}$ (86)
where $C_{1}=G(1+2x_{sup})$, $C_{2}=\frac{1}{K}+\log\big{(}\frac{T-1+K}{K}\big{)}$, $G$ is defined in (81) and $x_{sup}\coloneqq\sup\|x\|_{\mathcal{N},\overline{E}}$ is the upper bound of the $\|\cdot\|_{\mathcal{P}}$ norm for all $x^{t}$.
@@
This choice $\eta_{t}$ satisfies $\alpha_{3}\eta_{t}^{2}\leq\alpha_{2}\eta_{t}$. Thus, by (D.1.3) we have
@@
where the term $R_{1}$ is identical to the unbiased case in Theorem 3 of Zhang et al. (2021) which leads to
@@
also, $R_{2}$ can be bounded by a logarithmic dependence of $T$
@@
Appendix E Uncertainty Set Support Function Estimators
E.1 Proof of Theorem 3
@@
E.2 Proof of Theorem 4
denote $\hat{\sigma}^{*}_{\mathcal{P}^{a}_{s}}(V)$ as the untruncated MLMC estimator obtained by running Algorithm 1 when setting $N_{\max}$ to infinity. From Wang et al. (2023b), under both TV uncertainty sets and Wasserstein uncertainty sets, we have $\hat{\sigma}^{*}_{\mathcal{P}^{a}_{s}}(V)$ as an unbiased estimator of ${\sigma}_{\mathcal{P}^{a}_{s}}(V)$. Thus,
@@ E.3 Proof of Lemma 2
Combining (107) with (E.3) we obtain the first part of (23).
@@
using (111) and the fact that $|f(\lambda)-g(\lambda)|\leq\epsilon\Rightarrow|\sup_{\lambda}f(\lambda)-\sup_{\lambda}g(\lambda)|\leq\epsilon$, we obtain the second part of (23).
E.4 Proof of Theorem 5
@@
Appendix F Convergence for Robust TD
F.1 Formal Statement of Theorem 6
The first half of Algorithm 2 (line 1 - line 7) can be treated as a special instance of the SA updates in (66) with the bias and variance of the i.i.d. noise term specified in Section 4. To facilitate deriving the bounds of the noise terms, we first analyze the bounds in terms of the $l_{\infty}$ norm, and then translate the bounds in terms of the $\|\cdot\|_{\mathcal{P}}$ semi-norm to obtain the final results.
We start with analyzing the bias and variance of $\hat{\mathbf{T}}_{g_{0}}(V_{t})$ for each $t$. Recall the definition of $\hat{\mathbf{T}}_{g_{0}}(V_{t})$ is as follows:
@@ F.1 Formal Statement of Theorem 6
$\mathbb{E}$ $\left[(\hat{\mathbf{T}}_{g_{0}}(V_{t})(s)-{\mathbf{T}}_{g_{0}}(V_{t})(s))^{2}\right]$
$=\left(\mathbb{E}\left[\hat{\mathbf{T}}_{g_{0}}(V_{t})(s)\right]-{\mathbf{T}}_{g_{0}}(V_{t})(s)\right)^{2}+\mathrm{Var}\left(\hat{\mathbf{T}}_{g_{0}}(V_{t})(s)\right)$
@@
$t^{p}_{\mathrm{mix}}<+\infty\quad\text{and}\quad\|V\|_{\mathrm{sp}}\leq 4t^{p}_{\mathrm{mix}}\leq 4t_{\mathrm{mix}}$ (120)
where we define $t_{\mathrm{mix}}\coloneqq\sup_{p\in\mathcal{P}}t^{p}_{\mathrm{mix}}$, then $t_{\mathrm{mix}}$ is also finite due to the compactness of $\mathcal{P}$. We now derive the bounds of biases and variances for the three types of uncertainty sets. Regarding contamination uncertainty sets, according to Lemma 12, $\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)$ is unbiased and has variance bounded by $\|V\|^{2}$. Thu, define $t_{\mathrm{mix}}$ according to Lemma 13 and combining the above result with Lemma 13, we obtain that $\hat{\mathbf{T}}_{g_{0}}(V_{t})$ is also unbiased and the variance satisfies
@@
$\mathbb{E}$ $\left[\left\lVert\hat{\mathbf{T}}_{g_{0}}(V_{t})-{\mathbf{T}}_{g_{0}}(V_{t})\right\rVert^{2}_{\infty}\right]$
$\leq\left(24(1+\frac{1}{\delta})2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}\right)^{2}+3\|V\|_{\mathrm{sp}}^{2}+144(1+\frac{1}{\delta})^{2}\|V\|_{\mathrm{sp}}^{2}N_{\max}$
@@
In order to translate the above bounds from the $l_{\infty}$ norm into the $\|\cdot\|_{\mathcal{P}}$ norm, recall that in line 7 of Algorithm 2, we chose an anchor state $s_{0}$ set $V_{t}(s_{0})=0$ for all $t$ to avoid ambiguity. We thus can draw the following relationship:
Lemma 9.
Let $x\in\mathbb{R}^{S}$ satisfy $x_{i}=0$ for some fixed index $i$. Then
$\|x\|_{\infty}\leq\|x\|_{\mathrm{sp}}=\max_{1\leq j\leq n}x_{j}-\min_{1\leq j\leq n}x_{j}\leq 2\|x\|_{\infty}.$
Moreover, since all semi-norms with the same kernel spaces are equivalent, there are constants $c_{\mathcal{P}},C_{\mathcal{P}}>0$ so that
$c_{\mathcal{P}}\|x\|_{\mathrm{sp}}\leq\|x\|_{\mathcal{P}}\leq\;C_{\mathcal{P}}\|x\|_{\mathrm{sp}}\quad\forall\,x\in\mathbb{R}^{n},$
then
$c_{\mathcal{P}}\|x\|_{\infty}\leq c_{\mathcal{P}}\|x\|_{\mathrm{sp}}\leq\|x\|_{\mathcal{P}}\leq C_{\mathcal{P}}\|x\|_{\mathrm{sp}}\leq 2C_{\mathcal{P}}\|x\|_{\infty}.$ (126)
Proof.
Since $x_{i}=0$, for every $j$ we have $-\|x\|_{\infty}\leq x_{j}\leq\|x\|_{\infty}$. Hence
$\max_{j}x_{j}\;\leq\;\|x\|_{\infty},\quad\min_{j}x_{j}\;\geq\;-\,\|x\|_{\infty},$
and so
$\|x\|_{\infty}=\max\{\max_{j}x_{j}-0,0-\min_{j}x_{j}\}\leq\|x\|_{\mathrm{sp}}=\max_{j}x_{j}-\min_{j}x_{j}\leq\|x\|_{\infty}-(-\|x\|_{\infty})=2\,\|x\|_{\infty}.$
Since $\|\cdot\|_{\mathrm{sp}}$ and $\|\cdot\|_{\mathcal{P}}$ both have the same kernel of $\{c\mathbf{e}:c\in\mathbb{R}\}$, by the equivalence of semi-norms, it follows that there exists $c_{\mathcal{P}}$ and $C_{\mathcal{P}}$ such that
$c_{\mathcal{P}}\|x\|_{\infty}\leq c_{\mathcal{P}}\|x\|_{\mathrm{sp}}\leq\|x\|_{\mathcal{P}}\leq C_{\mathcal{P}}\|x\|_{\mathrm{sp}}\leq 2C_{\mathcal{P}}\|x\|_{\infty}$
as claimed. ∎
With the relationship established above, line 1 - line 7 of Algorithm 2 can be formally treated as a special instance of the SA updates in (66) with $B=0$. We now provide the bias and variance of the i.i.d. noise for the different uncertainty sets discussed using Lemma 9 and the estimation bounds in (121)-(F.1). For contamination uncertainty sets, we have
$\varepsilon^{\text{Cont}}_{\text{bias}}=0\quad\text{and}\quad A^{\text{Cont}}=32C_{\mathcal{P}}^{2}t^{2}_{\mathrm{mix}}$ (127)
@@
$\varepsilon^{\text{TV}}_{\text{bias}}=48C_{\mathcal{P}}(1+\frac{1}{\delta})2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}=\operatorname*{\mathcal{O}}\left(2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}\right)$ (128)
and
$A^{\text{TV}}$ $=2C_{\mathcal{P}}^{2}\left(24(1+\frac{1}{\delta})2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}\right)^{2}+96C_{\mathcal{P}}^{2}t_{\mathrm{mix}}^{2}+4608C_{\mathcal{P}}^{2}(1+\frac{1}{\delta})^{2}t_{\mathrm{mix}}^{2}N_{\max}$
$=\operatorname*{\mathcal{O}}\left(t_{\mathrm{mix}}^{2}N_{\max}\right)$ (129)
@@
$\varepsilon^{\text{Wass}}_{\text{bias}}=48C_{\mathcal{P}}\cdot 2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}=\operatorname*{\mathcal{O}}\left(2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}\right)$ (130)
and
$A^{\text{Wass}}$ $=2C_{\mathcal{P}}^{2}\left(24\cdot 2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}\right)^{2}+96C_{\mathcal{P}}^{2}t_{\mathrm{mix}}^{2}+4608C_{\mathcal{P}}^{2}t_{\mathrm{mix}}^{2}N_{\max}$
$=\operatorname*{\mathcal{O}}\left(t_{\mathrm{mix}}^{2}N_{\max}\right)$ (131)
@@
$\mathbb{E}\Bigl{[}\|V_{T}-V^{*}\|^{2}_{\infty}\Bigr{]}\leq\frac{4K^{2}c_{u}C_{\mathcal{P}}^{2}}{(T+K)^{2}c_{l}c_{\mathcal{P}}^{2}}\|V_{0}-V^{*}\|^{2}_{\infty}+\frac{8A^{\mathrm{Cont}}\alpha_{4}c_{u}}{(T+K)\alpha_{2}^{2}c_{\mathcal{P}}^{2}}=\operatorname*{\mathcal{O}}\left(\frac{1}{T^{2}}+\frac{t^{2}_{\mathrm{mix}}}{T(1-\gamma)^{2}}\right)$ (132)
@@
$\mathbb{E}\Bigl{[}\|V_{T}-V^{*}\|^{2}_{\infty}\Bigr{]}$ $\leq\frac{4K^{2}c_{u}C_{\mathcal{P}}^{2}}{(T+K)^{2}c_{l}c_{\mathcal{P}}^{2}}\|V_{0}-V^{*}\|^{2}_{\infty}+\frac{8A^{\mathrm{TV}}\alpha_{4}c_{u}}{(T+K)\alpha_{2}^{2}c_{\mathcal{P}}^{2}}+\frac{c_{u}C_{3}C_{2}\varepsilon^{\mathrm{TV}}_{\mathrm{bias}}}{\alpha_{2}c_{\mathcal{P}}^{2}}$ (133)
@@
$\mathbb{E}\Bigl{[}\|V_{T}-V^{*}\|^{2}_{\infty}\Bigr{]}$ $\leq\frac{4K^{2}c_{u}C_{\mathcal{P}}^{2}}{(T+K)^{2}c_{l}c_{\mathcal{P}}^{2}}\|V_{0}-V^{*}\|^{2}_{\infty}+\frac{8A^{\mathrm{Wass}}\alpha_{4}c_{u}}{(T+K)\alpha_{2}^{2}c_{\mathcal{P}}^{2}}+\frac{c_{u}C_{3}C_{2}\varepsilon^{\mathrm{Wass}}_{\mathrm{bias}}}{\alpha_{2}c_{\mathcal{P}}^{2}}$ (135)
@@
where the $\varepsilon$ and $A$ terms are defined in (127)-(F.1), $C_{2}=\frac{1}{K}+\log\big{(}\frac{T-1+K}{K}\big{)}$, $C_{3}=G(1+8C_{\mathcal{P}}t_{\mathrm{mix}})$, $\gamma$ is defined in (14), $c_{u},c_{l}$ are defined in (67), $\rho_{2}$ is defined in (70), $G$ is defined in (81), and $C_{\mathcal{P}},c_{\mathcal{P}}$ are defined in Lemma 9.